OpenStudy (anonymous):

If Mr. Cruz will increase his usual rate by 10 kilometers per hour, it will take him one hour shorter to cover a distance of 200 kilometers. What is his usual speed in driving?

OpenStudy (whpalmer4):

Let V be his usual rate, and T the usual time. V*T = 200 km We also know that if he speeds up by 10 km/h, the time goes down by an hour for the same distance. (V+10 km/h)(T-1 hr) = 200 km You should be able to solve that by substitution.

OpenStudy (whpalmer4):

You'll get a quadratic equation with two roots, but only one will make sense here.

OpenStudy (whpalmer4):

What's your answer?

OpenStudy (anonymous):

I still don't get it.

OpenStudy (ja1):

okay i just saw this question so what dont you get?

OpenStudy (anonymous):

what equation I can make out of the problem.

OpenStudy (ja1):

So the questions asks for an equation?

OpenStudy (anonymous):

no but i would need it to get the answer. I don't know how to get the answer.

OpenStudy (ja1):

well just one question what grade of math is this?

OpenStudy (anonymous):

what do you mean?

OpenStudy (ja1):

ike is it 6th grade algebra 7th 8th etc

OpenStudy (anonymous):

9th grade

OpenStudy (ja1):

basically just take this (V+10 km/h)(T-1 hr) = 200 km and substitute the variables. Or is that what you can't do?

OpenStudy (anonymous):

yes, it is.

OpenStudy (ja1):

well this is a bit out of my hands Im still in 8th but try bumping your question someone might see it.

OpenStudy (anonymous):

okay.. Thanks!

OpenStudy (whpalmer4):

\[(V + 10)(T-1) = 200\] \[VT - V + 10T -10 = 200\] But we know that VT = 200 — that's the usual time and speed. \[200 - V + 10T -10 = 200\]Subtract 200 from both sides \[10T - V -10 =0\]To solve this, we use VT=200 and solve it for V [\V = 200/T\] \[10T - (200/T) - 10 = 0\]Multiply both sides by T to eliminate fraction \[10T^2 - 200 - 10T = 0\] Now use the quadratic formula with a = 10, b = -10, c = -200 \[T = \frac{-b\pm\sqrt{b^2-4ac}}{2a} = \frac{-(-10)\pm\sqrt{(-10)^2-4(10)(-200)}}{20}\] \[T=\frac{10\pm\sqrt{100-(-8000)}}{20} = \frac{10\pm\sqrt{8100}}{20} = \frac{10\pm 90}{20} = 5, -40\] The only solution that makes sense here is T = 5, because he can't very well be taking -40 hours to drive somewhere. Now you know T, find his original velocity from V*T = 200. Then find the new velocity and time from (V + 10) and (T - 1) and make sure the solution works!